OpenStudy (anonymous):

5x + 8 − 3x = −10 Is the answer -9 or 9 ? I'm not sure because I got both..

OpenStudy (anonymous):

Aye, I may be of assistance

OpenStudy (vera_ewing):

5x + 8 -3x = -10 5x -3x + 8 - 8 = -10 - 8 2x = -18 x = -9

OpenStudy (anonymous):

But first I must know.. Where are the mayan key stones?

OpenStudy (vera_ewing):

So x = -9. Does that make sense?